a sum of money amounts to 2100 in a year and 2205 in two years . find the rate of interest and sum.



1
Expert's answer
2021-09-14T06:10:04-0400

Year 1 "=" 2100

Year 2 = 2205

"=" ?

"=" ?

Difference "=" 2205 "-" 2100

"=" 105

"=" p "\times" R "\times" T/100

105"=" "\frac {2100\times R\times1}{100}"

R"=" 5%

Part b

Let the principal for first year be x

So we'll use this formula;

"=" P "\times" R"\times" T/100

Whereby

"\frac{5x}{100}" "=" "\frac{x}{20}"

Amount for first year "=" x"+\frac{x}{20}" "=\frac{21x}{20}"

For second year p "=" "\frac{21x}{20}"

R= 5%p.a T = 1 year

So we solve for this;

"=" "[\frac{21x}{20}\times5\times1 ]/" 100

"=" "\frac{21x}{4}/" 100

From the equation above we'll get

"\frac{21x}{4}" /100

=2100

400"\times" "\frac{21x}{400}=" 2100"\times" 400

"\frac{21x}{21}" "=" "\frac{840000}{21}"

X"=" 40000

So the original sum of money is

40,000